find two consecutive even integers such that twice the first plus twice the second is 116
Answer by stanbon(75887)

find two consecutive even integers such that twice the first plus twice the second is 116
----
1st: 2x
2nd: 2x+2
----
Equation:
2(2x)+2(2x+2) = 116
-----
4x + 4x+4 = 116
---
8x = 112
---
x = 14
---
1st: 2x = 28
2nd: 2x+2 = 30
===================
Cheers,
Stan H.